# LLM-Based Multi-Agent Systems for Software Engineering: Vision and the Road Ahead

### v2

@@ Abstract.
Integrating Large Language Models (LLMs) into autonomous agents marks a significant shift in the research landscape by offering cognitive abilities competitive to human planning and reasoning. This paper explores the transformative potential of integrating Large Language Models into Multi-Agent (LMA) systems for addressing complex challenges in software engineering. LLMs enhance agents’ cognitive abilities, enabling more robust collaboration, autonomous problem-solving, and scalable project management. We review the current application of LMA systems in software engineering and discuss how LMA systems can improve software development through cross-examination, error reduction, and task automation. Additionally, the paper outlines the emerging challenges and presents a research agenda aimed at advancing LMA system capabilities in software engineering, offering key research questions to guide future exploration.
@@ 1. Introduction
Nevertheless, the application of singular LLM-based agents encounters limitations, since real-world problems often span multiple domains, requiring expertise from various fields. In response to this challenge, developing LLM-Based Multi-Agent (LMA) systems represents a pivotal evolution, aiming to boost performance via synergistic collaboration. An LMA system harnesses the strengths of multiple specialized agents, each with unique skills and responsibilities. These agents work in concert towards a common goal, engaging in collaborative activities like debate and discussion. These collaborative mechanisms have been proven to be instrumental in encouraging divergent thinking (Liang et al., 2023), enhancing factuality and reasoning (Du et al., 2023), and ensuring thorough validation (Wu et al., 2023). As a result, LMA systems hold promise in addressing a wide range of complicated real-world scenarios across various sectors (Horton, 2023; Wang et al., 2023e, a), such as software engineering (Lo, 2023; Qian et al., 2024c; Li et al., 2024; Hong et al., 2023).
@@
Existing research has illuminated the critical roles of these collaborative agents in advancing toward the era of Software Engineering 2.0 (Lo, 2023). LMA systems are expected to significantly speed up software development and drive innovation, marking a major shift in software engineering practices. This article aims to delve deeper into the roles of LMA systems in shaping the future of software engineering, spotlighting potential applications, emerging challenges, and the vital collaboration between academia and industry to harness these advancements fully. Furthermore, we present a comprehensive research agenda, which enumerates a series of research questions to direct subsequent research efforts.
@@ 3. Literature Review
In this section, we provide a review of recent studies on LMA systems in software engineering. We categorize these applications according to the various stages of the software development lifecycle (SDLC), focusing on key areas such as requirements engineering, code generation, quality assurance, and software maintenance. Additionally, we highlight papers that explore the use of LMA systems for end-to-end software development, which encompasses multiple phases of the SDLC, rather than focusing on isolated stages.
@@ 3.1. Requirements Engineering
Elicitron (Ataei et al., 2024) is an LMA framework that focuses specifically on the elicitation stage. It utilizes LLM-based agents to represent a diverse array of simulated users. These agents engage in simulated product interactions, providing insights into user needs by articulating their actions, observations, and challenges. MARE (Jin et al., 2024) is another LMA framework that covers multiple phases of requirements engineering, including elicitation, modeling, verification, and specification. It employs five distinct agents performing nine different tasks to help generate high-quality requirements models and specifications.
@@ 3.2. Code Generation
Code generation is an essential task in software development, where source code is produced from high-level specifications, often written in natural language. In this subsection, we review the recent innovative LMA frameworks that enhance the process of code generation.
A prominent design approach in LMA systems involves role specialization and cross-validation (e.g., peer review) among agents (Li et al., 2023; Liu et al., 2023; Sun et al., 2024b; Olausson et al., 2023; Dong et al., [n. d.]). In this method, agents are assigned distinct roles, such as programmers or testers, and collaborate iteratively to enhance code quality. INTERVENOR (Wang et al., 2024b) pairs a Code Learner with a Code Teacher to iteratively improve code and fix bugs using compiler error feedback. Similarly, Self-repair (Olausson et al., 2023) involves a code agent generating programs and a feedback agent providing execution error explanations to iteratively improve code. AutoCoder (Lei et al., 2024) simulates coding tasks with a questioner and a programmer. Both agents cooperate and verify the execution through unit tests. SoA (Ishibashi and Nishimura, 2024) proposes to organize agents based on problem complexity. PairCoder (Zhang et al., 2024a) incorporates a Navigator agent for high-level planning and a Driver agent for specific implementation. The Navigator is responsible for proposing promising solution plans, selecting the current optimal plan, and directing the next iteration round based on execution feedback. The Driver follows the guidance of Navigator to undertake initial code generation, code testing, and refinement. MapCoder (Islam et al., 2024) replicates the problem-solving process in competitive programming through four specialized LLM agents. The Retrieval Agent retrieves examples of similar problems, guiding the solution process. The Planning Agent then creates a structured solution plan, assigning confidence scores to each approach. The Coding Agent follows the plan to generate and test code, while the Debugging Agent continues to correct errors by comparing the code with the initial plan.
CodeS (Zan et al., 2024) converts natural language requirements into fully functional code repositories using three specialized agents. The RepoSketcher defines the repository structure, the FileSketcher generates blueprints for each file, and the SketchFiller completes the code for each function. Cai et al. (Cai et al., 2023) present a framework to enhance human-LLM interaction through visual programming techniques. It allows users to design workflows for complex tasks using low-code interfaces, giving them more control over LLM outputs. The framework includes two agents: a Planning agent, which creates workflows, and an Executing agent, which follows the user-edited workflows. CodexGraph (Liu et al., 2024b) improves code generation, code understanding, and other software engineering tasks by navigating large code repositories. It builds graph databases using static analysis, extracting code symbols and their relationships. These symbols, represented as nodes, and their interactions, like inheritance and containment, form the edges of the graph. CodexGraph enables LLM-based agents to interact with these graph databases by generating and executing graph queries.
@@ 3.3. Software Quality Assurance
Testing. Fuzz4All (Xia et al., 2024) generates testing input for software systems across multiple programming languages. In this framework, a distillation agent reduces user input, while a generation agent creates and mutates inputs. AXNav (Taeb et al., 2024) is designed to automate accessibility testing. It interprets natural language test instructions and executes accessibility tests, such as VoiceOver, on iOS devices. AXNav includes a planner agent, an action agent, and an evaluation agent. WhiteFox (Yang et al., 2023a) is a fuzzing framework that tests compiler optimizations. It uses two LLM-based agents: one extracts requirements from source code, and the other generates test programs. Additionally, LMA systems are employed for tasks such as penetration testing (Deng et al., 2023), user acceptance testing (Wang et al., 2024d), and GUI testing (Yoon et al., 2024).
Vulnerability Detection. GPTLens (Hu et al., 2023) is an LMA framework for detecting vulnerabilities in smart contracts. The system includes LLM-based agents acting as auditors, each independently identifying vulnerabilities. A critic agent then reviews and ranks these vulnerabilities, filtering out false positives and prioritizing the most critical ones. MuCoLD (Mao et al., 2024) assigns roles like tester and developer to evaluate code. Through discussions and iterative assessments, the agents reach a consensus on vulnerability classification.
@@ 3.4. Software Maintenance
Debugging. Debugging involves identifying, locating, and resolving software bugs. Several frameworks, including MASAI (Arora et al., 2024), MarsCode (Liu et al., 2024a), AutoSD (Kang et al., 2023), and others (Tao et al., 2024; Ma et al., 2024; Chen et al., 2024), follow a structured process consisting of stages like bug reproduction, fault localization, patch generation, and validation. Specialized agents are typically responsible for each stage. FixAgent (Lee et al., 2024) includes a debugging agent and a program repair agent that work together to iteratively fix code by analyzing both errors and repairs. The system refines fault localization by incorporating repair feedback. The agents also articulate their thought process, improving context-aware debugging. The MASTER framework (Yang et al., 2024) employs three specialized agents. The Code Quizzer generates quiz-like questions from buggy code, the Learner proposes solutions, and the Teacher reviews and refines the Learner’s responses. AutoCodeOver (Zhang et al., 2024c) uses an agent for fault localization via spectrum-based methods, collaborating with others to refine patches using program representations like abstract syntax trees. SpecRover (Ruan et al., 2024) extends AutoCodeOver by improving program fixes through iterative searches and specification analysis, based on inferred code intent. ACFIX (Zhang et al., 2024b) targets access control vulnerabilities in smart contracts, focusing on Role-Based Access Control. It mines common RBAC patterns from over 344,000 contracts to guide agents in generating patches. DEI (Zhang et al., 2024f) resolves GitHub issues by using a meta-policy to select the best solution, integrating and re-ranking patches generated by different agents for improved issue resolution. RepoUnderstander (Ma et al., 2024) constructs a knowledge graph for a full software repository and uses Monte Carlo tree search to assist in understanding complex dependencies.
@@ 3.5. End-to-end Software Development
End-to-end software development involves managing the entire software development lifecycle, from gathering initial requirements to delivering the final product. In the context of LMA systems, each phase typically is handled by specialized agents which perform tasks suited to their expertise. The ultimate goal is to create fully autonomous systems capable of managing the entire development cycle. These systems often draw inspiration from established software development practices, such as Agile (Cohen et al., 2004), Scrum (Schwaber, 1997), and Waterfall (Petersen et al., 2009).
Pioneering works, such as ChatDev (Qian et al., 2024c) and MetaGPT (Hong et al., 2023), adopt a Waterfall model to automate the software development. ChatDev assigns agents as programmers, reviewers, and testers to handle design, coding, testing, and documentation. MetaGPT enhances the workflow by incorporating Standardized Operating Procedures (SOPs) (Belbin and Brown, 2022), with agents taking on roles like product managers and QA engineers. Recent advancements (Zhang et al., 2024d; Dong et al., 2023; Du et al., 2024) have further built on these designs, pushing the boundaries of fully autonomous development. Self-collaboration (Dong et al., 2023) presents a framework for improving code generation using multiple instances of ChatGPT. The approach simulates teamwork by assigning specific roles (analyst, coder, and tester) to distinct instances of ChatGPT. CTC (Du et al., 2024) involves multiple teams of LLM-based agents working together to generate high-quality software. Each team independently proposes solutions, and cross-communication allows them to refine their outputs by eliminating suboptimal proposals.
AgileCoder (Nguyen et al., 2024) and AgileGen (Zhang et al., 2024e) apply Agile methodologies in software development. AgileCoder (Nguyen et al., 2024) mirrors Agile’s role-based collaboration, assigning tasks like Product Manager and Developer to work in sprints. It also features a Dynamic Code Graph Generator, which updates code dependencies in real time, supporting context-aware generation and refinement. AgileGen (Zhang et al., 2024e) combines human-AI collaboration with Agile principles, using Gherkin for user acceptance criteria, ensuring alignment between requirements and code, and reducing user effort.
TGen (Mathews and Nagappan, 2024) explores Test-Driven Development (TDD), where test cases guide GPT-4 to generate correct code. LCG (Lin et al., 2024a) emulates various software process models, including Waterfall, TDD, and Scrum.
Additionally, several works (Rasheed et al., 2024b; Lin et al., 2024b; Qian et al., 2024b; Sami et al., 2024) adopt iterative and adaptive methodologies without explicitly detailing their models. For instance, ToP (Lin et al., 2024b) utilizes process models to guide agents in software development, which are refined through heuristic algorithms and process mining (Van Der Aalst, 2012). Qian et al. (Qian et al., 2024b) propose an iterative experience refinement framework that enables agents to continuously adapt to software development by leveraging prior experiences. This framework supports ongoing learning through the acquisition, utilization, and selective elimination of experiences. Co-Learning (Qian et al., 2024a) fosters collaboration through co-tracking, co-memorizing, and co-reasoning modules among agents. GameGPT (Chen et al., 2023b) automates and streamlines game development, employing various roles like game designers and code reviewers to enhance decision-making in game creation.

### v3

# LLM-Based Multi-Agent Systems for Software Engineering: Literature Review, Vision and the Road Ahead

@@ Abstract.
Integrating Large Language Models (LLMs) into autonomous agents marks a significant shift in the research landscape by offering cognitive abilities that are competitive with human planning and reasoning. This paper explores the transformative potential of integrating Large Language Models into Multi-Agent (LMA) systems for addressing complex challenges in software engineering (SE). By leveraging the collaborative and specialized abilities of multiple agents, LMA systems enable autonomous problem-solving, improve robustness, and provide scalable solutions for managing the complexity of real-world software projects. In this paper, we conduct a systematic review of recent primary studies to map the current landscape of LMA applications across various stages of the software development lifecycle (SDLC). To illustrate current capabilities and limitations, we perform two case studies to demonstrate the effectiveness of state-of-the-art LMA frameworks. Additionally, we identify critical research gaps and propose a comprehensive research agenda focused on enhancing individual agent capabilities and optimizing agent synergy. Our work outlines a forward-looking vision for developing fully autonomous, scalable, and trustworthy LMA systems, laying the foundation for the evolution of Software Engineering 2.0.
@@ 1. Introduction
Nevertheless, the application of singular LLM-based agents encounters limitations, since real-world problems often span multiple domains, requiring expertise from various fields. In response to this challenge, developing LLM-Based Multi-Agent (LMA) systems represents a pivotal evolution, aiming to boost performance via synergistic collaboration. An LMA system harnesses the strengths of multiple specialized agents, each with unique skills and responsibilities. These agents work in concert towards a common goal, engaging in collaborative activities like debate and discussion. These collaborative mechanisms have been proven to be instrumental in encouraging divergent thinking (Liang et al., 2023), enhancing factuality and reasoning (Du et al., 2023), and ensuring thorough validation (Wu et al., 2023b). As a result, LMA systems hold promise in addressing a wide range of complicated real-world scenarios across various sectors (Horton, 2023; Wang et al., 2023e, a), such as software engineering (Lo, 2023; Qian et al., 2024c; Li et al., 2024a; Hong et al., 2023).
@@
Existing research has illuminated the critical roles of these collaborative agents in advancing toward the era of Software Engineering 2.0 (Lo, 2023). LMA systems are expected to significantly speed up software development, drive innovation, and transform the current software engineering practices. This article aims to delve deeper into the roles of LMA systems in shaping the future of software engineering. It spotlights the current progress, emerging challenges, and the road ahead. We provide a systematic review of LMA applications in SE, complemented by two case studies that assess current LMA systems’ capabilities and limitations. From this analysis, we identify key research gaps and propose a comprehensive agenda structured in two phases: (1) enhancing individual agent capabilities and (2) optimizing agent collaboration and synergy. This roadmap aims to guide the development of autonomous, scalable, and trustworthy LMA systems, paving the way for the next generation of software engineering
To summarize, this study makes the following key contributions:
  • We conduct a systematic review of 71 recent primary studies on the application of LMA systems in software engineering.
  • We perform two case studies to illustrate the current capabilities and limitations of LMA systems.
  • We identify the key research gaps and propose a structured research agenda that outlines potential future directions and opportunities to advance LMA systems for software engineering tasks.
@@ 3. Literature Review
In this section, we review recent studies on LMA systems in software engineering, organizing these applications across various stages of the software development lifecycle, including requirements engineering, code generation, quality assurance, and software maintenance. We also examine studies on LMA systems for end-to-end software development, covering multiple SDLC phases rather than isolated stages.
Search Strategy: We conduct a keyword-based search on the DBLP publication database (DBLP Computer Science Bibliography, 2024) to match paper titles. DBLP is a widely used resource in software engineering surveys (Chen et al., 2020; Zhang et al., 2018; Chen et al., 2024b), which indexes over 7.5 million publications across 1,800 journals and 6,700 academic conferences in computer science.
Our search included two sets of keywords: one set targeting LLM-based Multi-Agent Systems (called [agent words]) and the other focusing on specific software engineering activities (called [SE words]). Papers may use variations of the same keyword. For example, the term “vulnerability” may appear as “vulnerable” or “vulnerabilities.” To address this, we use truncated terms like “vulnerab” to capture all related forms. For LMA systems, we used keywords: “Agent” OR “LLM” OR “Large Language Model” OR “Collaborat”. To ensure comprehensive coverage of SE activities, we incorporated phase-specific keywords for each stage of the SDLC into our search queries:
  1. (1)
    Requirements Engineering: requirement, specification, stakeholder
  2. (2)
    Code Generation: software, code, coding, program
  3. (3)
    Quality Assurance: bug, fault, defect, fuzz, test, vulnerab, verificat, validat
  4. (4)
    Maintenance: debug, repair, review, refactor, patch, maintenanc
We focus on four key phases of the SDLC: requirements engineering, code generation, quality assurance, and software maintenance. For each phase, the relevant SE keywords are combined using the OR operator to capture all variations. The final search query for each SDLC phase follows the format: [agent words] AND [SE words].
Following the guide of previous work (Meline, 2006; Zhou et al., 2024; Van Dinter et al., 2021), we design the following inclusion and exclusion criteria. In the first phase, we filtered out short papers (exclusion criterion 1) and removed duplicates (exclusion criterion 2). In the second phase, we manually screened each paper’s venue, title, and abstract, excluding items such as books, keynote speeches, panel summaries, technical reports, theses, tool demonstrations, editorials, literature reviews, and surveys (exclusion criteria 3 and 4). In the third phase, we conducted a full-text review to further refine relevant studies. Following Section 2.3, we exclude papers that do not describe LMA systems (exclusion criterion 5). Papers that rely solely on LLMs using non-agent-based methods or single-agent approaches are excluded. Further, we focused on LMA systems powered by LLMs with strong planning capabilities, such as ChatGPT and LLaMA, excluding models like CodeBERT and GraphCodeBERT. Since the release of ChatGPT is in November 2022, we limited our review to papers published after this date (exclusion criterion 6). Furthermore, we excluded papers unrelated to software engineering (exclusion criterion 7) and those that mention LMA systems only in discussions or as future work, without presenting experimental results (exclusion criterion 8). After the third phase, we identified 41 primary studies directly relevant to our research focus. The search process is conducted on November 14th, 2024.
  • The paper must be written in English.
  • The paper must have an accessible full text.
  • The paper must adopt LMA techniques to solve software engineering-related tasks.
  • The paper has less than 5 pages.
  • Duplicate papers or similar studies authored by the same authors.
  • Books, keynote records, panel summaries, technical reports, theses, tool demos papers, editorials
  • The paper is a literature review or survey.
  • The paper does not utilize LMA systems, e.g., using a single LLM agent.
  • The paper is published before November 2022 (the release date of ChatGPT).
  • The paper does not involve software engineering related tasks.
  • The paper lacks experimental results and mentions LMA systems only in future work or discussions.
Snowballing Search To expand our review, we conducted both backward and forward snowballing (Wohlin, 2014) on the relevant papers identified in previous steps. This process involved examining the references cited by the relevant studies as well as publications that have cited these studies. We repeated the snowballing process until reaching a transitive closure fixed point, where no new relevant papers were found, resulting in an additional 30 papers identified.
@@ 3.1. Requirements Engineering
Elicitron (Ataei et al., 2024) is an LMA framework that focuses specifically on the elicitation stage. It utilizes LLM-based agents to represent a diverse array of simulated users. These agents engage in simulated product interactions, providing insights into user needs by articulating their actions, observations, and challenges. MARE (Jin et al., 2024) is an LMA framework that covers multiple phases of requirements engineering, including elicitation, modeling, verification, and specification. It employs five distinct agents, i.e., stakeholder, collector, modeler, checker, and documenter, performing nine actions to help generate high-quality requirements models and specifications. Sami et al. (Sami et al., 2024b) propose another LMA framework to generate, evaluate, and prioritize user stories through a collaborative process involving four agents: product owner, developer, quality assurance (QA), and manager. The produce owner generates user stories and initiates prioritization. The QA agent assesses story quality and identifies risks, while the developer prioritizes based on technical feasibility. Finally, the manager synthesizes these inputs and finalizes prioritization after discussions with all agents
@@ 3.2. Code Generation
Code generation (Herrington, 2003; Budinsky et al., 1996) has consistently been a longstanding focus of software engineering research, aiming to automate coding tasks to boost productivity and minimize human error.
A prominent multi-agent setup for code generation typically on role specialization and iterative feedback loops to optimize collaboration among agents. We summarize the common roles identified in the literature, including the Orchestrator, Programmer, Reviewer, Tester, and Information Retriever.
The Orchestrator acts as the central coordinator, managing high-level planning and ensuring smooth task execution across all agents. Its responsibilities include defining high-level strategic goals, breaking them into actionable sub-tasks, delegating these tasks to the appropriate agents, monitoring progress, and ensuring that workflows align with overall project objectives (Li et al., 2024c; Zhang et al., 2024a; Ishibashi and Nishimura, 2024; Zan et al., 2024; Cai et al., 2023; Phan et al., 2024; Li et al., 2024a; Josifoski et al., 2023). For instance, PairCoder (Zhang et al., 2024a) features a Navigator agent that interprets natural language descriptions to create high-level plans outlining solutions and key implementation steps. The Driver agent then follows these plans to handle code generation and refinement. The Self-Organized Agents (SoA) framework (Ishibashi and Nishimura, 2024) employs a hierarchical design, with Mother agents managing high-level abstractions and delegating subtasks to specialized Child agents. In CODES (Zan et al., 2024), the Orchestrator role is performed by the RepoSketcher, which converts high-level natural language requirements into a repository sketch. This sketch outlines the project structure, including directories, files, and inter-file dependencies. The RepoSketcher then delegates tasks to the FileSketcher and SketchFiller, ensuring the efficient and seamless creation of a complete, functional code repository.
During the implementation phase, the process typically begins with the Programmer, who is responsible for writing the initial version of the code. Once the initial code is produced, roles like the Reviewer and Tester step in to evaluate it, providing constructive feedback on quality, functionality, and adherence to requirements. This feedback initiates an iterative cycle, where the Programmer refines the code or the Debugger resolves identified issues, ensuring that the final code meets the desired standards and performs as expected (Mathews and Nagappan, 2024; Wang et al., 2024b; Olausson et al., 2023; Chen et al., 2023c; Le et al., 2024; Liu et al., 2023; Lei et al., 2024a; Lin et al., 2024a; Dong et al., 2023). For example, INTERVENOR (Wang et al., 2024b) pairs a Code Learner with a Code Teacher. The Code Learner generates the initial code and then compiles it to evaluate its correctness. If issues are identified, the Code Teacher analyzes the bug reports and the buggy code, subsequently providing repair instructions to address the errors. Self-repair (Olausson et al., 2023) and TGen (Mathews and Nagappan, 2024) refine code by utilizing feedback obtained from running pre-defined test cases.
When predefined test cases are unavailable, the Tester can generate a variety of test cases, ranging from common scenarios to edge cases. These tests help uncover subtle issues that might otherwise go unnoticed and provide actionable feedback to guide subsequent refinement iterations (Huang et al., 2023; Shinn et al., 2024; Ishibashi and Nishimura, 2024; Hu et al., 2024).
Some frameworks employ the Information Retriever to gather relevant information to assist code generation. For instance, Agent4PLC (Liu et al., 2024d) and MapCoder (Islam et al., 2024) incorporate a Retrieval Agent tasked with sourcing examples of similar problems and extracting related knowledge. This agent provides essential contextual information and references tailored to the user’s input, ensuring that solutions are well-informed and adhere to domain-specific best practices. Similarly, CodexGraph (Liu et al., 2024b) employs a translation agent to facilitate interaction with graph databases, which are built using static analysis to extract code symbols and their relationships. By converting user queries into graph query language, this agent enables precise and structured information retrieval, enhancing the capability of LLM-based agents to navigate and utilize code repositories effectively.
Agent Forest (Li et al., 2024d) adopts a different paradigm instead of role specialization. Instead, it utilizes a sampling-and-voting framework, where multiple agents independently generate candidate outputs. Each output is then evaluated based on its similarity to the others, with a cumulative similarity score calculated for each. The output with the highest score—indicating the greatest consensus among the agents—is selected as the final solution.
@@ 3.3. Software Quality Assurance
Testing. Fuzz4All (Xia et al., 2024) generates testing input for software systems across multiple programming languages. In this framework, a distillation agent reduces user input while a generation agent creates and mutates inputs. AXNav (Taeb et al., 2024) is designed to automate accessibility testing. It interprets natural language test instructions and executes accessibility tests, such as VoiceOver, on iOS devices. AXNav includes a planner agent, an action agent, and an evaluation agent. WhiteFox (Yang et al., 2023a) is a fuzzing framework that tests compiler optimizations. It uses two LLM-based agents: one extracts requirements from source code, and the other generates test programs. Additionally, LMA systems are employed for tasks such as penetration testing (Deng et al., 2023), user acceptance testing (Wang et al., 2024e), and GUI testing (Yoon et al., 2024).
Vulnerability Detection. GPTLens (Hu et al., 2023) is an LMA framework for detecting vulnerabilities in smart contracts. The system includes LLM-based agents acting as auditors, each independently identifying vulnerabilities. A critic agent then reviews and ranks these vulnerabilities, filtering out false positives and prioritizing the most critical ones. MuCoLD (Mao et al., 2024) assigns roles like tester and developer to evaluate code. Through discussions and iterative assessments, the agents reach a consensus on vulnerability classification. Widyasari et al. (Widyasari et al., 2024) introduces a cross-validation technique, where multiple LLM’s answer is validated against each other.
@@ 3.4. Software Maintenance
Debugging. Debugging involves identifying, locating, and resolving software bugs. Several frameworks, including MASAI (Arora et al., 2024), MarsCode (Liu et al., 2024a), AutoSD (Kang et al., 2023), and others (Tao et al., 2024; Ma et al., 2024; Chen et al., 2024a; Lei et al., 2024b), follow a structured process consisting of stages like bug reproduction, fault localization, patch generation, and validation. Specialized agents are typically responsible for each stage. FixAgent (Lee et al., 2024) includes a debugging agent and a program repair agent that work together to iteratively fix code by analyzing both errors and repairs. The system refines fault localization by incorporating repair feedback. The agents also articulate their thought processes, improving context-aware debugging. The MASTER framework (Yang et al., 2024) employs three specialized agents. The Code Quizzer generates quiz-like questions from buggy code, the Learner proposes solutions, and the Teacher reviews and refines the Learner’s responses. AutoCodeOver (Zhang et al., 2024c) uses an agent for fault localization via spectrum-based methods, collaborating with others to refine patches using program representations like abstract syntax trees. SpecRover (Ruan et al., 2024) extends AutoCodeOver by improving program fixes through iterative searches and specification analysis based on inferred code intent. ACFIX (Zhang et al., 2024b) targets access control vulnerabilities in smart contracts, focusing on Role-Based Access Control. It mines common RBAC patterns from over 344,000 contracts to guide agents in generating patches. DEI (Zhang et al., 2024f) resolves GitHub issues by using a meta-policy to select the best solution, integrating and re-ranking patches generated by different agents for improved issue resolution. SWE-Search (Antoniades et al., 2024) consists of three agents: the SWE-Agent for adaptive exploration, the Value Agent paired with a Monte Carlo tree search module for iterative feedback and utility estimation, and the Discriminator Agent for collaborative decision-making through debate. RepoUnderstander (Ma et al., 2024) constructs a knowledge graph for a full software repository and also uses Monte Carlo tree search to assist in understanding complex dependencies.
@@
Test Case Maintenance. Lemner et al. (Lemner et al., 2024) propose two multi-agent architectures to predict which test cases need maintenance after source code changes. These agents perform tasks including summarizing code changes, identifying maintenance triggers, and localizing relevant test cases.
@@ 3.5. End-to-end Software Development
End-to-end software development encompasses the entire process of creating a software product. While conventional code generation is often limited to producing isolated components such as functions, classes, or modules, end-to-end development starts from high-level software requirements and progresses through design, implementation, testing, and ultimately delivering a fully functional and ready-to-use product.
In practice, developers and stakeholders typically adopt established software process models to guide collaboration, such as Agile (Cohen et al., 2004) and Waterfall (Petersen et al., 2009). Similarly, the design of LMA systems for end-to-end software development draws inspiration from these software process models. The development process is organized into distinct phases, such as requirements gathering, software design, implementation, and testing. Each phase is managed by specialized agents with domain expertise.
Figure 1. Multi-Agent Systems in Software Development: Waterfall vs. Agile Models
It is important to note that works such as FlowGen (Lin et al., 2024a) and Self-Collaboration (Dong et al., 2023) emulate various software process models. However, their experiments focus on generating code segments rather than delivering fully developed software products. As a result, in this paper, these approaches are not considered to be designed for true end-to-end software development.
Several works (Qian et al., 2024c; Hong et al., 2023; Zhang et al., 2024d; Du et al., 2024; Zan et al., 2024; Sami et al., 2024a; Rasheed et al., 2024b; Holt et al., 2023) adopt the Waterfall model to automate software development. The Waterfall model used in these multi-agent methods organizes the software development process into distinct, sequential phases, where each stage must be completed before proceeding to the next. The primary phases typically include Requirement Analysis, Architecture Design, Code Development, Testing, and Maintenance. For instance, in MetaGPT (Hong et al., 2023), the Product Manager agent thoroughly analyzes user requirements. The Architect agent then transforms these requirements into detailed system design components. Subsequently, the Engineer implements the specified classes and functions as outlined in the design. Finally, the Quality Assurance Engineer creates and executes test cases to ensure rigorous code quality standards are met. These approaches emphasize a linear and sequential design process, ensuring structured progression and clear accountability at each stage.
AgileCoder (Nguyen et al., 2024) and AgileGen (Zhang et al., 2024e) adopt Agile process models for software development, emphasizing iterative development by breaking complex tasks into small, manageable increments. AgileCoder (Nguyen et al., 2024) assigns Agile roles such as Product Manager and Scrum Master to facilitate sprint-based collaboration and development cycles. AgileGen enhances Agile practices with human-AI collaboration, integrating close user involvement to ensure alignment between requirements and generated code. A notable feature of AgileGen is its use of the Gherkin language to create testable requirements, bridging the gap between user needs and code implementation.
While most methods rely on predefined roles and fixed workflows for software development, a few work (Wang et al., 2024d; Li et al., 2023; Lin et al., 2024b) investing in dynamic process models. Think-on-Process (ToP) (Lin et al., 2024b) introduces a dynamic process generation framework. Since software development processes can vary significantly depending on project requirements, ToP moves beyond the limitations of static, one-size-fits-all workflows to enable more flexible and efficient development practices. Given a software requirement, this framework leverages LLMs to create tailored process instances based on their knowledge of software development. These instances act as blueprints to guide the architecture of the LMA system, adapting to the specific and diverse needs of different projects. Similarly, in MegaAgent (Wang et al., 2024d), agent roles and tasks are not predefined but are generated and planned dynamically based on project requirements. Both ToP and MegaAgent highlight the shift from rigid, static workflows to dynamic, adaptive systems. These frameworks promise more efficient, flexible, and context-aware software development practices, aligning processes with project-specific requirements and complexities.
Additionally, instead of focusing on the process model, several works (Qian et al., 2024a, b) explore leveraging experiences from past software projects to enhance new software development efforts. Co-Learning (Qian et al., 2024a) enhances agents’ software development abilities by utilizing insights gathered from historical communications. This framework fosters cooperative learning between two agent roles—instructor and assistant—by extracting and applying heuristics from their task execution histories. Building on this, Qian et al. (Qian et al., 2024b) propose an iterative experience refinement (IER) framework that enables agents to continuously adapt by acquiring, utilizing, and selectively refining experiences from previous tasks, improving agents’ effectiveness and collaboration in dynamic software development scenarios.
4. Case Study
To demonstrate the practical effectiveness of LMA systems, we conduct two case studies. Specifically, we utilize the state-of-the-art LMA framework, ChatDev (Qian et al., 2024c), to autonomously develop two classic games: Snake and Tetris. ChatDev structures the software development process into three phases: designing, coding, and testing. ChatDev employs specialized roles, including CEO, CTO, programmer, reviewer, and tester. ChatDev’s agents are powered by GPT-3.5-turbo [^4] https://platform.openai.com/docs/models/gp#gpt-3-5-turbo. The temperature setting controls the randomness and creativity of the GPT-3.5’s responses. Following the original ChatDev setting, we set the temperature of GPT-3.5-turbo as 0.2.
4.1. Snake Game
For the Snake game, we provide the following prompt to ChatDev to generate the game:
Snake Game Prompt “Design and implement a grid-based snake game displayed on the screen. Initialize the snake with a defined starting position, length, and direction. Enable continuous movement controlled by arrow keys. Introduce food that spawns randomly on the grid, ensuring it does not overlap with the snake. Trigger snake growth when food is consumed, adding a new segment to its body. Implement a game-over condition for boundary or body collisions, displaying a message and providing a restart option. Include a scoring system displayed in the user interface, along with clear instructions.
While the first attempt to generate the Snake game was unsuccessful, we resubmitted the same prompt to ChatDev, and the second attempt successfully produced a playable version. ChatDev also generated a detailed manual that included information on dependencies, step-by-step instructions for running the game, and an overview of its features. Figure LABEL:fig:snake displays the graphical user interfaces (GUIs) of the generated Snake game, showing the starting state, in-game state, and game-over state. The development process was consistently efficient, taking an average of 76 seconds and costing \$0.019. Upon playing the game, we confirmed that it fulfilled all the requirements outlined in the prompt.
4.2. Tetris Game
We present the following prompt to ChatDev to guide the generation of the Tetris game:
Tetris Game Prompt Design and implement a Tetris game. Start with a randomly chosen piece dropping from the top. Allow players to control the tetromino using arrow keys for movement (left, right, down) and rotation. Enable automatic downward movement with an adjustable speed. Handle collisions with the boundaries and existing pieces, locking the tetromino in place when it cannot move further. Check for complete rows after each placement and remove them. End the game if new tetrominoes cannot spawn due to a full board, displaying a game over message.
Figure 2. Screen shots of the Tetris Game generated by ChatDev.
During development, ChatDev faced challenges in producing functional gameplay across the first nine attempts. Notice that the same prompt was used for each run. On the tenth attempt, ChatDev successfully produced a Tetris game that met most of the prompt requirements, as shown in Figure 3. The figure illustrates the game’s key states: the starting state, in-game states, and game-over state. However, the game still lacks the core functionality to remove completed rows, as demonstrated in the third subplot of Figure 3. Overall, the development process remained efficient, with an average time of 70 seconds and a cost of \$0.020 per attempt.
Figure 3. Screen shots of the Tetris Game generated by ChatDev.
Summary of Findings. From our case studies, current LMA systems demonstrate strong performance in reasonably complex tasks like developing a Snake game. The generated Snake game meets all requirements in the prompt within just a few iterations. The process was efficient and cost-effective, with an average completion time of 76 seconds and a cost of \$0.019 per attempt. These results emphasize the suitability of LMA systems for moderately complex software engineering tasks. However, when tasked with more complex challenges like developing a Tetris game, ChatDev successfully generates a playable Tetris game only by the tenth attempt. The game still lacks the core functionality, i.e., removing completed rows. This highlights the limitations of current LMA systems in handling more complex tasks that require deeper logical reasoning and abstraction. Nevertheless, development remains efficient and cost-effective, averaging 70 seconds and \$0.020 per run, making the system a promising tool for rapid prototyping.